If 5 is added to the sum of two consecutive integers, the result is three less than twice the larger integer. Find the integers.

I keep getting 8

Answer by venugopalramana(3286)

If 5 is added to the sum of two consecutive integers, the result is three less than twice the larger integer. Find the integers.
LET 2 CONSECUTIVE INTEGERS BE ...N,N+1
SUM OF THE 2 NUMBERS PLUS 5 =N+N+1+5=2N+6
TWICE LARGER INTEGER =2(N+1)=2N+2
3 LESS THAN THIS =2N+2-3=2N-1
SO 2N+6=2N-1.....THIS HAS NO SOLUTION..CHECK THE PROBLEM..